When x number of units are sold, the price of each unit (in dollars) is given by P=-x/2+75. Find the unit price when the following quantities are sold: 2,7,9,11.
Answer by tutorcecilia(2152)

P=-x/2+75 [plug-in each value of x and solve:
.
P=-(2)/2+75 [let x=2]
P=74
.
P=-(7)/2+75 [let x=7]
P=71.5
.
P=-(9)/2+75 [let x=9]
P=70.5
.
P=-(11)/2+75 [let x=11]
P=69.6